CRY+VIS Interface Meeting
Date: 2021/12/10 13:00-
Acceptance Check Meeting on Suspension.
Date: 2021/12/10 14:00-15:00
Participants: Takahashi, Fabian, Ishizaki, Tanaka, Terrence, Aso, Ushiba, Uchiyama
- Repaired the thermometer cable in IX.
- Cleaned the cross tube of EXV with the vacuum cleaner.
- Investigated the BF V2 LVDT in EX. The noise level may be increased due to the large DC position.
- Repaired the LVDT driver for the GAS filters in IX. The original main card was noisy.
- Recovered the yaw FR in F0 in EY. The helix-coil spring was too loose.
- Investigated the actuator of the BF damper in EY. The cable from the Dsub connector was touching the bracket for the secondary coil.
- Post O4, we are investigating whether SiC can be used as a base plate material to the payload increasing of GAS filters.
- Terrence fitted SR2 TFs with mathematical models and install them in filters banks DAMP/OLDAMP F10.
- Then he used a method, coded in his Python library Kontrol, to design the control filters. Read the klog for details. The main resonances are well supressed.
- In SR2 he added comb filters to LVDTs filter banks to get rid of the 60 Hz peak and its harmonics.
- Terrence reported that the polarity of the BS F1 LVDT actuation unit was reversed.
- We confirmed the adapter box reverses the polarity of the LVDT.
We should change the polarity of the primary coil within the box: this requires exchanging places of pins 2 and 7.
- Fabian acquired geophone data from the DAQ in units of counts in order to compare with previous measurements and theoretical estimations.
- The IPs were swinging free on a day of good weather.
- IN SR2, the new measurement is consistent with the three-point correlation analysis did by Fujii-kun, rather than with the theoretical estimate.
- Terrence reported that the BS IP was very soft, with a resonant frequency in IP-L of 47 mHz. I double-checked and reviewed previous measurements.
I corroborated his assessment:50 and 82 mHz for L and T respectively (5 mHz resolution).
Last time we adjusted the resonant frequencies (21-06-2021) they were higher: 63 and 92 mHz.
- The temperature during the ballast mass adjustment and the recent assessments were very similar, so I cannot blame the temperature for this.
- I don't know what could have produced this change.
We should remove some ballast mass next week, keeping 47 mHz is risky.
- What are we going to do about the BS F1 LVDT?
- I think we should keep the adapter box. It improves the SNR. We should exchange pins 2 and 7. I guess we aim to go beyond the current prototype.
- Connecting the chamber to the rack GND does not make the wandering noise disappear. We should assess how it will affect the performance of the control.
- We aim to calibrate the LVDT the week after the next one becuase of the X arm commissioning work next week. However, next week we'd like to test the sensor mounting hardware if possible.
- Type-A team: IMMT suspension. Pump down of the IXV and EXV.
- Type-B team: IP tuning in BS, F1 LVDT noise in BS, SRM control
- Task list for authorship in VIS
- The author list making process was revised at the F2F meeting in August. As the first step, the author list committee required each subsystem chief to make the task list in each subsystem.
The task list (Draft) should be submitted to SEO by 15th Dec.
- The tasks listed are not just plans but the tasks carried out within 2021. If you have any tasks on VIS not listed but contributed by you, please let me know.
(Date in Kamioka)
The week from 12/13
- Takahashi: 12/13-17
- Sato: 12/13-17
- Hirata: 12/13-17
The week from 12/20
- Takahashi: 12/20-24
- Sato: 12/20-24
- Hirata: 12/20-24